OpenStudy (anonymous):

Algebra 2 please help

OpenStudy (anonymous):

Solve 10^x + 75 = 150. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

Okay! hold on a sec:)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

i got 7.5 then i dont know what to do

OpenStudy (anonymous):

A. x ≈ 1.63 B. x ≈ 1.56 C. x ≈ 1.75 D. x ≈ 1.88 i am pretty sure its A

OpenStudy (anonymous):

Okay, let me try something real quick.

OpenStudy (anonymous):

It's D! just look at this:

OpenStudy (anonymous):

ohhh ok i get it thank you :) <3

OpenStudy (anonymous):

No prob! if you need help with anything just feel free to ask!

OpenStudy (anonymous):

you can just take the log of each... \[10^x=75 \Rightarrow x=\log 75\] use your calculator to find log 75 = 1.87506...